# Are Biological Systems More Intelligent Than Artificial Intelligence?

### v5

@@ Abstract
Are biological self-organising systems more ‘intelligent’ than artificial intelligence (AI)? If so, why? I explore this through a mathematical lens which frames intelligence in terms of adaptability. I model systems as stacks of abstraction layers (Stack Theory) and compare them by how they delegate agentic control down their stacks, illustrating with examples of computational, biological, human military, governmental and economic systems. Contemporary AI rests on a static, human-engineered stack in which lower layers are static during deployment. Put provocatively, static stacks resemble inflexible bureaucracies, adapting only top-down. Biological stacks are more ‘intelligent’ because they delegate adaptation. Formally, I prove a theorem (The Law of the Stack) showing adaptability in higher layers requires sufficient adaptability in lower layers. Generalising bio-electric explanations of cancer as isolation from collective informational structures, I explore how cancer-like failures occur in non-biological systems when delegation is inadequate. This helps explain how to build more robust systems, by delegating control like the military doctrine of mission command. It also provides a design perspective on hybrid agents (e.g. organoids, systems involving both humans and AI): hybrid creation is a boundary-condition design problem in which human-imposed constraints prune low-level policy spaces to yield desired collective behaviour while preserving collective identity.
@@ 1 Introduction
Biological systems self-organise; their parts and the environment interact to produce spatiotemporal order [ashby1947, seeley2002, rosas2018, bell2007, kelso1997, Friston2010, tognoli2014]. They pursue homeostatic and reproductive goals. Some proactively learn [kukushkin2024], storing information about the past in order to anticipate future needs. To make accurate predictions in complex environments, a system must learn cause and effect [pearl2018, merker2005, barron2016]. I consider a system’s ‘intelligence’ to be a measure of how efficiently it does this [bennett2024a, chollet2019, bennettmaruyama2022b, bennett2023c] [^1] Throughout, “intelligence” is used in an efficiency-of-adaptation sense (sample and energy efficiency). For a comparison with other common definitions, see Electronic Supplementary Materials (ESM) Note S1.. In this view intelligence is a long-term adaptation hard-wired into the organism, that enables short-term adaptation during its lifetime [bennett2025thesis].
Artificial intelligence (AI) is the study of adaptive machines, and machine learning (ML) is the study of machines that learn [russell2005ai, bishop2006, sutton2018, goertzel2014]. Contemporary approaches to machine learning are often inefficient [bennettmaruyama2022a]. They can require vast quantities of data and energy to learn tasks a human can pick up after one or two examples. Even organoids [^2] By organoid I mean a collective of cells grown in a lab, placed in a computational interface and life support system, and then ‘trained’ to communicate with conventional computing technology. appear more ‘intelligent’ in this sense [khajehnejad2024]. For example, an organoid outperformed conventional reinforcement learning systems at the game Pong, adapting with greater sample efficiency. That organoid was a highly tuned hybrid system whose behaviour was shaped by human-imposed boundary conditions. This is similar to how an acellular slime mold can be made to ‘solve’ shortest-path problems [nakagaki2000]. These sorts of results are highly dependent upon the human imposed boundary conditions, but they raise an interesting question. Are biological self-organising systems more intelligent than artificial intelligence? If so, why?
Some say the key difference between biological systems and contemporary artificial intelligence is that the former are active inference systems, while the latter are typically not [parr2024, Friston2010, friston2013]. It is true that active inference can improve efficiency [ororbia2022, ororbia2023, ororbia2023b]. Active inference has been applied in the context of the ARC-AGI benchmark [chollet2019], which to some extent tests sample efficiency. Active inference systems performed well (for example the 2024 ARC-AGI submission by Jack Cole), but still underperform humans. Hence there may be more to the picture [^3] From the perspective of this paper, Active Inference/FEP is primarily a theory of how an agent updates beliefs and selects actions (e.g. via variational free energy minimisation), whereas this paper is more about where and at what level of abstraction adaptation is realised in a system’s stack. Hence I treat the approaches as potentially complementary. An extended comparison and related work is provided in ESM Note S6..
Others seem to suggest biological systems are not more intelligent [^4] In the sense of sample and energy efficiency of adaptation., and the difference is primarily one of inductive bias [chollet2019]. This inductive bias is hard-coded foreknowledge of the environment, which makes it easier to learn and adapt to that environment. To whatever extent this is true, it substitutes the question ‘what is intelligence?’ for ‘what is the inductive bias?’.
To reiterate, I want to know why biological self-organisation seems more adaptive [^5] Sample and energy efficient. than contemporary AI. Is there something else biological systems do that allows them to be more ‘intelligent’? Could there be more to the picture than active inference and an inductive bias?
I suggest there is. Computers are organised into layers of abstraction, each layer being the behaviour of the layer below. For example, software runs on hardware. If software is making decisions, then it controls the hardware top-down. In dominant contemporary AI deployments, adaptation is concentrated at a comparatively high level of abstraction (e.g. model parameters and software-level state), while the lower computational stack (instruction set, microarchitecture, hardware) is treated as fixed during deployment. [^6] This is not a claim about a fundamental limitation of computational systems: reconfigurable or self-modifying substrates (e.g. FPGAs, adaptive firmware, agentic materials, or bio-hybrid substrates) can in principle shift adaptation downward. The claim here is about the dominant architectures and workflows in contemporary AI. In contrast, biological self-organisation uses the multiscale competency architecture (MCA). Adaptation takes place at every scale, from cells to organs to organisms [fields2020]. It is ‘scale-free’. The same dynamics are observed at different scales. Control is delegated, in the sense that individual cells pursue goals [^7] As opposed to in a computer where the parts are usually designed to passively await instruction.. Cells are a “soft” material that has agency [man2019, ball2023]. A cell has a perceptual field. A cell can co-operate with other cells in this perceptual field. This allows them to form multicellular networks that have a larger perceptual field than any individual cell, to maintain coherent higher level “identity” like that of an organism [levin2024]. Each cell has its own goals and is intelligent in the sense of being able to reach those goals by various means [james1890, ciaunica2023]. However each cell can also be part of a collective of cells that in turn pursues its own collective goals. Smaller collectives can form larger collectives, and the pattern repeats at larger scales. Collectives are nested [ciaunica2021, ciaunica2023b]. Each “thinks” in the enactive sense, interacting with its environment [thompson2007, varela2016, rolla2021]. Biological collectives can exhibit high level goal directed behaviour far beyond the capability of the parts of which they’re formed [reid2011].
@@
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent [^8] In the specific efficiency-based sense defined above. than contemporary, predominantly fixed-stack implementations of AI [^9] This is not a claim about all possible future AI. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack.. This is because biological self-organisation delegates adaptation to lower levels of abstraction [^10] For readers who want the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2.. Formally, previous work showed adaptability in embodied systems can be measured by the ‘weakness’ of the constraints they embody [bennett2025d, bennett2024a]. Building on this, I prove a theorem (The Law of the Stack) that the ability to embody weak constraints at higher levels of abstraction depends on the ability to embody weak constraints at lower levels. This result shows how we can build more robust and efficient hybrid and cyberphysical systems, laying a foundation for future empirical research. Its applications extend beyond artificial intelligence, for example to human organisational and economic systems.
@@ 2 Abstraction, Distribution and Delegation
Conventional computing systems are organised into stacks of abstraction layers. More generally, I will characterise systems along three interacting dimensions: abstraction (how layers relate), distribution (how work is spread within a layer), and delegation (the level of abstraction at which adaptation or control resides). How efficiently a system adapts depends on the interaction between these dimensions.
In proposing Stack Theory [^11] Initially proposed in my PhD thesis [bennett2025thesis]. [bennett2025thesis], I merely extend this idea to systems in general. I’ll begin illustrating what I mean by abstraction, delegation and distribution; dimensions along which one may plot any system. For intuition, I will give examples of limit cases for computational, biological, human organisational and economic systems.
@@ 2.1 Abstraction
In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA [^12] This is the software library that allows for the large scale utilisation of GPUs (a type of hardware originally developed for games), for machine learning and artificial intelligence., which runs on a GPU [^13] A graphical processing unit (GPU) is a type of hardware that allows for massively parallel vector processing. It was developed for the purpose of graphics processing in videogames, and later repurposed for machine learning..
@@
These are all examples of abstraction layers. A programming language is a sort of abstraction layer. For example Python is an abstraction layer over C, in the sense that it is interpreted by a program written in the less abstract C programming language. C in turn depends on a program written in Assembly language, which runs on machine code which is interpreted by hardware. Each layer of language is written in the language of the layer below. This structure allows programmers to work more efficiently, within the confines of a standardised and simplified representation of the system. However, these abstraction layers are incomplete, simplified pictures of the system. They “abstract away” unnecessary details. Most importantly, they are embodied in the layer below. Each language is a subset of the behaviour that the layer below can exhibit. For example, a subset of the possible behaviour a given computer can exhibit would correspond to running Python code, and a broader subset of that computer’s behaviour would necessarily correspond to running C code. Hence each layer in this stack is a subset of the behaviours of the layers below [^14] A higher-level abstraction is not uniquely determined by its lower-level implementation: the same abstract interface can be realised by multiple implementations. The claim here is that higher-level behaviour must be realised by lower-level behaviour, and that different implementations can differ in resource cost (time, energy, information), which matters for adaptability.. This reductionist framing does not deny that useful abstractions can be robust to lower-level variation; invariants can be treated as higher-level task constraints. Even the physical hardware itself can be framed as an abstraction layer over the environment, being a subset of the behaviour of that environment in which the hardware exists [bennett2025thesis].
Think of a stack of abstraction layers as a directed graph. Highly abstract behaviours are a consequence of ‘weaker’, more generalisable goal directed behaviour at lower levels of abstraction. Framed this way, everything is a stack of abstraction layers [^15] This is not an analogical or metaphorical application of computational frames to others systems, as in Bratton’s re-framing of geopolitics [bratton2016]. Rather, this is a more general claim about how all systems work.. For example a collective of cells can enact an organ, and a system of organs can enact an organism [levin2024] which enacts even more abstract behaviours like language or tools. Each of these is an ‘abstraction layer’ in a biological stack [bennett2024a].
@@ 2.2 Distribution
In the case of biological systems lower levels of abstraction are formed of smaller or simpler parts [delgado1997], which is to say the work is distributed among parts at a level of abstraction. Because of this, looking at a lower level of abstraction is like ‘zooming in’ to look at a smaller scale. Conversely, rather than looking at low levels of abstraction like cells I can shift the level of abstraction to higher in the stack, and look at individual humans and how they form collectives at higher levels of abstraction. For example a population of soldiers can be a squad, and if squads co-operate they can be a platoon. Each layer above is the behaviour of a layer below, and the parts get smaller when I go down.
@@
In each layer work is distributed among members of the population. If level of abstraction is one dimension along which a system may be plotted, then distribution of work is another. In a computer I can distribute work by splitting it up into sub-tasks and allocating it to many computers.
However, unlike biological systems a lower level of abstraction in a computer is not necessarily made up of smaller parts that form a collective. Looking at a lower level of abstraction is not like ‘zooming in’. I can run software on a solitary computer, or a cluster of computers. For example, Python code is run by an interpreter, which is a C program that in turn is built on Assembly, which is built on machine code that has a fixed interpretation physically hard-wired in the instruction set architecture of the CPU. The scale of the smallest part at each layer of abstraction is the same. Think of one Python program running on one interpreter, on one computer. This illustrates how level of abstraction can vary without affecting the scale or distribution.
@@ 2.3 Delegated Control
The aforementioned stack on which Python relies is not a collective. It is not made up of parts capable of functioning independently. They do not have agency. This means control of the system resides at another level of abstraction. If I am programming a computer, then control resides at the level of abstraction at which I am programming. If I am programming in Python, then adaptation occurs at the Python level of abstraction [^16] This is a statement about where learning/adaptation is typically implemented in conventional software stacks. Lower layers can of course carry mutable state (data, caches, firmware state), but are usually not designed to be autonomously goal-directed in the way MCA components are..
Running a program in a distributed system does not always affect this. Distributing work is not the same thing as delegating control or agency. Work can be distributed among many GPUs, but if I remove a GPU from the computer it will not then pursue goals independently. It will sit, inert until once again integrated into a larger system. It requires no incentives. Control is imposed top-down, from higher to lower levels of abstraction. The parts at lower levels have no agency.
@@
Conversely in a biological system each cell is independently capable of goal directed behaviour [ball2023], assuming it remains alive. Control in a complex biological system begins bottom-up, before top-down. Control then propagates in both directions [^17] In the sense that members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with.. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs. Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
@@
A good example to understand the limit cases for delegation is an economy. In an absolutely centrally planned economy, control is concentrated at the highest level of abstraction. One might think of the people in the economy as parts at lower level of abstraction. They do work but have no control. Those people could adapt, but central planner’s edicts constrain them. The opposite extreme would be to have control delegated entirely to the lowest level of abstraction, in anarchy unconstrained by any laws.
The level of abstraction is often related to the granularity at which the system is represented. For example if the level of abstraction at which I view an economic system is humans, then I don’t worry about how work is distributed or control is delegated within those humans. I just concern myself with what systems can be formed of humans, like economies or armies (see fig 1). Of course, there are limits on what sort of physical forms can store or process information in service of a goal [sole2024], and this may limit how much I can distribute work or delegate control. Delegating to a lower level of abstraction often requires distributing work among parts at smaller scales [^18] For example, delegating control to the level of individual humans would mean those humans control their own work potential, but not that of others, or else we’d just be delegating to collectives within the human population.. However as the static abstraction layers in a computer illustrate, a lower level of abstraction is not always more granular, and thus work is not always more distributed (see fig 2).
@@
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [bennett2025thesis]).
@@
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [bennett2025thesis]).

### v6

@@ Abstract
Are biological self-organising systems more “intelligent” than artificial intelligence (AI)? If so, why? I address this question using a mathematical framework that defines intelligence in terms of adaptability. Systems are modelled as stacks of abstraction layers (Stack Theory) and compared by how effectively they delegate agentic control down their stacks. I illustrate this using computational, biological, military, governmental, and economic systems. Contemporary AI typically relies on static, human-engineered stacks whose lower layers are fixed during deployment. Put provocatively, such systems resemble inflexible bureaucracies that adapt only top-down. Biological systems are more intelligent because they delegate adaptation. Formally, I prove a theorem (The Law of the Stack) showing that adaptability at higher layers is bottlenecked by adaptability at lower layers. I further show that, under standard viability assumptions, maximising adaptability is equivalent to minimising variational free energy, implying that delegation is necessary for free-energy minimisation. Generalising bioelectric accounts of cancer as isolation from collective informational structures, I analyse cancer-like failure modes in non-biological systems when delegation is inadequate. This yields design principles for building robust systems via delegated control, and reframes hybrid agents (e.g. organoids or human–AI systems) as weak boundary-condition design problems in which constraints shape low-level policy spaces while preserving collective identity.
@@ 1 Introduction
Organoids are living systems whose behaviour is sculpted by engineered constraints. Human brain cells embedded in a computational interface have been trained to play Pong, achieving remarkable sample efficiency compared to a contemporary reinforcement learning agent [53]. Similarly, Physarum slime mould is normally an indiscriminate expander. However it reliably computes shortest paths when placed inside a maze [62]. These behaviours are neither programmed nor accidental. They arise because biological agency is coupled to carefully designed boundary conditions. Such systems are hybrid agents, jointly produced by bottom-up cellular dynamics and top-down environmental structure [82].
Biological systems self-organise across scales. Their parts and environment interact to produce spatiotemporal order [4, 78, 76, 9, 51, 39, 92]. Many systems proactively learn, storing information about the past to anticipate future needs [54]. To adapt in complex environments, a system must infer cause and effect [72, 60, 7]. Throughout, I use ‘intelligence’ in the specific sense of the sample and energy efficiency of adaptation [16, 26, 11, 14] [^1] Mainstream AI typically treats intelligence as performance across tasks. My conclusions concern efficiency of acquiring said performance. Accordingly, claims about “AI” should be read as applying to dominant contemporary ML training and deployment stacks, rather than all possible systems. See ESM Note S1..
Two common explanations for biological efficiency are active inference under the Free Energy Principle (FEP) [39, 40, 71] and powerful inductive biases [26]. These address, respectively, how agents update beliefs and act, and what structure is pre-specified. This paper isolates a complementary structural question, namely where in a system is adaptation permitted?
Computers are organised into layers of abstraction. In dominant AI deployments, adaptation is concentrated at high software layers (model parameters and software-level state) while lower layers (instruction set, microarchitecture, hardware) are treated as fixed during deployment. [^2] Lower layers can carry mutable state (data, caches, firmware), and self-modifying substrates exist; the claim here is about the prevailing engineering pattern in mainstream ML systems. In contrast, biological multiscale competency architectures (MCA) delegate goal-directed control across scales, from cells to organs to organisms [34, 59]. I model this difference with Stack Theory, formalising systems as stacks of abstraction layers. Delegation then measures how far adaptive control reaches down the stack.
@@
The somewhat provocative claim this paper will make is that biological self-organising systems are more intelligent (in the efficiency-of-adaptation sense) than contemporary, predominantly fixed-stack implementations of AI. This is not a claim about all possible future AI. One aim of the paper is to identify design principles for building more adaptable hybrid and cyberphysical systems by shifting where adaptation occurs in the stack. This is because biological self-organisation delegates adaptation to lower levels of abstraction (for the generalisation of “levels of abstraction” beyond computing, see ESM Supplementary Note S2). Formally, I measure adaptability by the weakness of the constraints a system embodies [20, 16]. I then prove The Law of the Stack (Theorem 5.1), which is that utility (adaptability) achievable one layer up is bounded by the weakness of the realised policy one layer down. Finally, I link this to the FEP. Under a simple viability prior, maximising weakness is equivalent to minimising variational free energy (see ESM “Viability, lived history, and variational free energy”), which makes delegation a prerequisite for sustained free energy minimisation across scales.
For clarity, the core contributions can be read as four claims:
Claim 1.1 (Delegation constrains adaptability).
In a multilayer architecture, increasing adaptability at an interface layer requires sufficiently weak policies at lower layers (formalised by Theorem 5.1).
Claim 1.2 (Delegation is necessary for free energy minimisation).
Under a simple viability prior, variational free energy minimisation reduces to keeping options open (maximising entropy over viable futures). The Law of the Stack then implies that abstract-layer free energy cannot be reduced without sufficient lower-layer slack, making delegation a structural prerequisite for sustained free energy minimisation.
Claim 1.3 (Insufficient delegation can induce cancer-analogue fragmentation).
Model a hybrid at some vocabulary $v$ as a distributed system with part-tasks $\alpha_{1},\dots,\alpha_{m}$ and a collective task $\omega$ whose collective-policy set is
$\Pi_{\omega}\;:=\;\bigcap_{j=1}^{m}\Pi_{\alpha_{j}}.$
(Thus the collective has an identity at $v$ iff $\Pi_{\omega}\neq\emptyset$.) Model top-down boundary conditions / insufficient delegation as additional constraints that restrict each part’s feasible policies to a subset $\widehat{\Pi}_{\alpha_{j}}\subseteq\Pi_{\alpha_{j}}$, inducing
$\widehat{\Pi}_{\omega}\;:=\;\bigcap_{j=1}^{m}\widehat{\Pi}_{\alpha_{j}}.$
If these restrictions render the system over-constrained at $v$ (i.e. $\widehat{\Pi}_{\omega}=\emptyset$), then no common policy exists at that level of abstraction. If at least one part remains locally viable (i.e. $\widehat{\Pi}_{\alpha_{j}}\neq\emptyset$ for some $j$), then any continued function must proceed via a proper subset of parts whose restricted policy sets still admit a non-empty intersection. I refer to this formal situation as “cancer-like” purely by analogy.
Claim 1.4 (Hybrid creation as maximally weak boundary-condition design).
Hybrid agents (e.g. organoids, soft robots, human–AI organisations) are best treated as boundary-condition design problems, which I formalise using Stack Theory. One must engineer constraints that prune low-level policy spaces just enough to realise target collective behaviour while preserving a non-empty intersection of component policy sets (collective identity). Formalised this way, the optimal choice of boundary conditions are the weakest sufficient constraints to achieve the desired hybrid functionality.
The remainder of the paper develops the intuition (Section 2), states the formal framework and theorem, and applies the result to hybrid agency and failure modes. The discussion of potential applications spans computational, biological, human governmental and economic systems.
@@ 2 Abstraction, Distribution and Delegation
Conventional computing systems are organised into stacks of abstraction layers. More generally, I characterise systems along three interacting dimensions. Abstraction describes how layers relate. Distribution describes how work is spread within a layer. Delegation is the level of abstraction at which adaptation or control resides. How efficiently a system adapts depends on the interaction between these dimensions.
In proposing Stack Theory [19], I extend this idea beyond computing. I illustrate abstraction, distribution, and delegation with limiting-case examples from computing, biology, and human organisations.
@@ 2.1 Abstraction
In a computer, a predictive model typically relies on a machine learning library, which runs on CUDA (a GPU programming platform), which runs on a graphics processing unit (GPU).
@@
Abstraction layers are interfaces that hide lower-level details. In computation, Python is interpreted by C, which runs on assembly and machine code on hardware. Each higher-layer behaviour must be realised by some lower-layer behaviour, but the same abstract interface can be realised by multiple implementations. This many-to-one mapping is where time, energy and information costs hide, so the abstraction layer through which an agent senses and acts constrains its adaptability. In Stack Theory, useful invariants are treated as higher-level task constraints even when micro-implementations vary. Even hardware itself can be framed as an abstraction layer over the local environment [19].
Think of a stack as a directed graph of constraints. Higher-level behaviours are implemented by and limited by lower-level behaviours. In biology, cells enact organs, organs enact organisms, and organisms enact still more abstract behaviours such as communication and tool use [59, 16].
@@ 2.2 Distribution
Distribution refers to how work is spread within a layer. In many biological systems, lower levels of abstraction are literally smaller parts [32], so “going down” is like zooming in. Cells form organs and organisms form organisms. For human organisations, soldiers form squads and squads form platoons.
@@
Distribution is distinct from abstraction, in that the same abstraction layer can be realised by one component or many.
In computation, lower abstraction does not necessarily mean smaller spatial parts. A single Python program is interpreted by a C program, which relies on assembly and machine code executed by the same physical CPU. Thus one can change abstraction without changing the scale or distribution of underlying components.
@@ 2.3 Delegated Control
Delegation is the level of abstraction at which adaptive control resides. In a typical computing stack, lower layers (instruction set, microarchitecture, hardware) implement fixed semantics rather than autonomous goal-directed behaviour. Adaptation therefore occurs where policy can change. This is usually at the software layer we program, even though lower layers can carry mutable state (data, caches, firmware).
Distribution is not delegation. A workload can be spread across many GPUs without granting any component independent control. Remove a GPU and it does not continue pursuing the system’s objectives, or any other for that matter.
@@
Biological collectives differ in that cells and tissues are autonomously goal-directed at their own scale [6]. Higher-level coherence arises when the viable policy sets of parts intersect, producing a collective policy (collective identity). Control can propagate bottom-up and top-down by which I mean collectives constrain their parts (e.g. via signals or incentives), but parts also shape collective behaviour. For illustrative purposes, consider how members of a collective are constrained by the collective, for example members of a human society are incentivised to obey laws. This is a form of top-down control. Bottom-up then becomes just whatever the individual humans do within those constraints, even though the system as a whole might have emerged entirely bottom-up from anarchy to begin with. Each part of a collective acts on its own, and its actions imply a ‘policy’. Each individual policy determines how that part interacts with the other parts, or more to the point their policies. These interactions cause the system as a whole to act, which implies a collective policy. If collective and individual policies align in service of homeostasis, then the system can persist. It has a collective identity or goal. That identity can constrain parts of the collective. For example, I might say the level of abstraction is cells, and then describe an organ as a collective policy of those cells. Likewise, I might look at the level of organs, and describe a human as a collective policy of those organs. Or I could describe the human as the collective policy of cells, and see the organs as a sort of middle management layer.
@@
Economies illustrate useful limit cases to understand delegation. One might think of the people in the economy as parts at lower level of abstraction. Central planning concentrates control at the top, while extreme decentralisation delegates control to the bottom. In the central planning limit case, the population does work but has no control. In the limit those people could adapt given local information unavailable to the central planner [44], but central planner’s edicts constrain them. In the opposite extreme there is no coordination.
Real systems occupy a continuum. Of course, there are limits on what sort of physical forms can store or process information in service of a goal [81], and this may limit how much I can distribute work or delegate control. Delegating control often coincides with shifting to finer granularity and greater distribution. However, as computers illustrate, abstraction can also change without changing the spatial scale of components (Figures 1 and 2).
@@
Figure 1: Control or agency and thus adaptation in a given system can take place at a particular level of abstraction. For example, I can program in Python, or I can go down a level and program in C. If I program in C it doesn’t matter whether I program one computer or fifty. Either way my control is at the same level of abstraction (figure from [19]).
@@
Figure 2: Work, resources and so on can be distributed across a system. Distribution is distinct from delegation of control. For example, both food stamps and universal basic income might distribute the same amount of resources to the same people within an economic system, but the former more tightly constrains the recipients in how they use those resources (figure from [19]).